# latexOverwriting file

### v4

@@ Abstract
The ensemble Kalman filter is widely used in applications because, for high dimensional filtering problems, it has a robustness that is not shared for example by the particle filter; in particular it does not suffer from weight collapse. However, there is no theory which quantifies its accuracy as an approximation of the true filtering distribution, except in the Gaussian setting. To address this issue we provide the first analysis of the accuracy of the ensemble Kalman filter beyond the Gaussian setting. We prove two types of results: the first type comprise a stability estimate controlling the error made by the ensemble Kalman filter in terms of the difference between the true filtering distribution and a nearby Gaussian; and the second type use this stability result to show that, in a neighbourhood of Gaussian problems, the ensemble Kalman filter makes a small error, in comparison with the true filtering distribution. Our analysis is developed for the mean field ensemble Kalman filter. We rewrite the update equations for this filter, and for the true filtering distribution, in terms of maps on probability measures. We introduce a weighted total variation metric to estimate the distance between the two filters and we prove various stability estimates for the maps defining the evolution of the two filters, in this metric. Using these stability estimates we prove results of the first and second type, in the weighted total variation metric. We also provide a generalization of these results to the Gaussian projected filter, which can be viewed as a mean field description of the unscented Kalman filter.
@@ 1.1 Context
This paper is concerned with the study of partially and noisily observed dynamical systems. Filtering refers to the sequential updating of the probability distribution of the state of the (possibly stochastic) dynamical system, given partial noisy observations [3, 43, 51]. The Kalman filter, introduced in 1960, provides an explicit solution to this problem in the setting of linear dynamical systems, linear observations and additive Gaussian noise [38]; the desired probability distribution is Gaussian and the Kalman filter provides explicit update formulae for the mean and covariance. The extended Kalman filter is a linearization-based methodology which was developed in the 1960s and 1970s to apply to situations beyond the linear-Gaussian setting [36]. It is, however, not practical in high dimensions, because of the need to compute, and sequentially update, large covariance matrices [29]. To address this issue, the ensemble Kalman filter was introduced in 1994 [25], using ensemble-based low-rank approximations of the covariances, and is a method well-adapted to high dimensional problems. The unscented Kalman filter, introduced in 1997, provides an alternative approach to nonlinear and non-Gaussian problems [37], using quadrature to approximate covariance matrices, and is well-adapted to problems of moderate dimension. The particle filter [20] is a provably convergent methodology for approximating the filtering distribution [11, 50]. However it does not scale well to high dimensional problems [4, 56]; this motivates the increasing adoption of ensemble Kalman methods.
@@
Over the last two decades ensemble Kalman filters have found widespread use in the geophysical sciences, are starting to be used in other application domains, and have been developed as a general purpose tool for solving inverse problems; for reviews of such developments see [8, 24, 26]. Despite widespread and growing adoption, theory quantifying the accuracy of the ensemble Kalman filter, in relation to the true filtering distribution, is limited. Two exceptions are the important contributions [44, 46] which concern accuracy in the large particle limit in the setting where the underlying filtering problem is Gaussian. However, there is no proof that the ensemble Kalman methodology can accurately approximate the desired filtering distribution beyond the Gaussian setting. Indeed, in general the methodology does not correctly reproduce the filtering distribution; for examples and analysis in the setting of filtering and Bayesian inverse problems see [1, 44] and [22] respectively. The aim of our work is to address this issue by proving accuracy of the ensemble Kalman filter beyond the Gaussian setting; specifically our analysis applies when the true filtering distribution is close to Gaussian, after appropriate lifting to the joint space of state and data. We perform the analysis for the mean field ensemble Kalman filter, focusing on quantifying the effect of the Gaussian approximation underlying the ensemble Kalman filter. We also study the Gaussian projected filter; this filter can be seen as a mean field version of the unscented Kalman filter. Both the ensemble and unscented Kalman mean field models are defined in [8].
@@
Although our study of the accuracy of ensemble Kalman filters, beyond the Gaussian setting, is new, there exists a growing body of literature analyzing ensemble Kalman methods from different perspectives. In the context of long-time behaviour see [39, 60, 13, 12]. For filter divergence see [30, 40]. For localization see [59, 58]. For high dimensions see [45]. For continuous-time and mean-field limits see [41, 42, 23]. In the context of inverse problems see [35, 54, 55, 34, 33].
@@ 1.3 Notation
where $\mathrm{KL}(\mu\|\nu)$ is the Kullback–Leibler (KL) divergence of $\mu$ from $\nu$, defined in (B.46). Note that $\mathsf{G}$ defines a nonlinear mapping. We refer to $\mathsf{G}$ as a projection because of its characterization as finding closest point to $\mu$ with respect to $\mathrm{KL}(\mu\|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}).$ Throughout this paper all probability measures have Lebesgue density, because of our assumptions concerning the noise structure in the dynamics model and the data acquisition model. Thus we abuse notation by using the same symbols for probability measures and their densities. For a probability measure $\mu\in\mathcal{P}(^{n})$, the notation $\mu(u)$ for $u\in^{n}$ refers to the Lebesgue density of $\mu$ evaluated at $u$, whereas $\mu[f]$ for a function $f\colon^{n}\to$ is a short-hand notation for $\int_{{}^{n}}f\mathrm{d}\mu$.
@@
Let $g\colon^{n}\to[1,\infty)$ be given by $g(v)=1+\lvert v\rvert^{2}$. We define the weighted total variation metric $d_{g}\colon\mathcal{P}(^{n})\times\mathcal{P}(^{n})\to$ by
@@
Metrics of this type have been employed previously in the literature. See for example the early reference [48], where a weighted total variation metric is used for stuyding ergodicity for Markov chains, and the reference [47], where a similar metric is employed in the context of perturbation of Markov chains. See also [32], where a direct proof of Harris’ ergodic theorem relying on appropriate weighted total variation norms is given, as well as [63, Theorem 6.15], which states that Wasserstein distances can be controlled by weighted total variation metrics. There are other dual metrics on probability measures commonly employed in the literature, such as the so-called maximum mean discrepancy, which is based on a dual formulation in a reproducing kernel Hilbert space; see e.g. [57].
@@
If $\mu_{1},\mu_{2}$ have Lebesgue densities $\rho_{1}$, $\rho_{2}$, then
@@
Remark 3.
Unlike the usual total variation distance, the weighted metric in Definition 1 enables control of the differences $\lvert\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\rvert$ and $\lVert\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\rVert$; several lemmas used to prove our main results rely on this control. More precisely, it is possible to show that for all $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ with finite second moments, it holds that
@@ 1.4.1 General Case
Here $\{u_{j}\}_{j\in\llbracket 0,J\rrbracket}$ is the unknown state, evolving in ${}^{d}$, and $\{y_{j}\}_{j\in\llbracket 0,J\rrbracket}$ are the observations, evolving in ${}^{K}$. We assume that the initial state is distributed according to a Gaussian distribution $u_{0}\sim\mathcal{N}(m_{0},\mathit{C}_{0})$ and that the following independence assumption is satisfied:
@@
The objective of probabilistic filtering is to sequentially estimate the conditional distribution of the unknown state, given the data, as new data arrives. The true filtering distribution $\mu_{j}$ is the conditional distribution of the state $u_{j}$ given a realization $Y_{j}^{\dagger}=\{y_{1}^{\dagger},\dotsc,y_{j}^{\dagger}\}$ of the data process up to step $j$. Data $Y_{j}$ may be thought of as arising from a realization of (1.4.1); but the case of model mis-specification, where $Y_{j}$ does not necessarily arise from (1.4.1), is also of interest.
It is well-known [43, 51] that the true filtering distribution evolves according to
@@
where $\mathsf{P}$ and $\mathsf{L}_{j}$ are maps on probability measures, referred to respectively as the prediction and analysis steps in the data assimilation community [3]. The operator $\mathsf{P}$ is linear and defined by the Markov kernel associated with the stochastic dynamics (1.2a). Its action on a probability measure with Lebesgue density $\mu$ reads
@@
The operator $\mathsf{L}_{j}$ is a nonlinear map which formalizes the incorporation of the new datum $y^{\dagger}_{j+1}$ using Bayes’ theorem. Its action reads
@@
The operator $\mathsf{L}_{j}$ effects a reweighting of the measure to which it applies, with more weight assigned to the state values that are consistent with the observation. It is convenient in this work to decompose the analysis map $\mathsf{L}_{j}$ into the composition $\mathsf{B}_{j}\mathsf{Q}$, where the operators $\mathsf{Q}\colon\mathcal{P}(^{d})\to\mathcal{P}(^{d}\times^{K})$ and $\mathsf{B}_{j}\colon\mathcal{P}(^{d}\times^{K})\to\mathcal{P}(^{d})$ are defined by
@@
and
@@ 1.4.2 Gaussian Case
Now consider the setting in which $\Psi(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})=M\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}$, and $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})=H\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}$ for matrices $\mathit{M}\in^{d\times d}$ and $\mathit{H}\in^{K\times d}.$ In this case the filtering distribution is Gaussian: $\mu_{j}=\mathcal{N}(m_{j},C_{j}).$ The Kalman filter [38] gives explicit evolution equations for the pair $(m_{j},C_{j})\in^{d}\times^{d\times d}.$ To write these down it is helpful to note that $\widehat{\mu}_{j+1}:={\mathsf{P}}\mu_{j}$ is also Gaussian: $\widehat{\mu}_{j+1}=\mathcal{N}(\widehat{m}_{j},\widehat{\mathit{C}}_{j}).$
@@
where $\{{y}^{\dagger}_{j}\}$ is the observation. Recall that this observation may be thought of as arising from a realization of (1.4.1); but, in the case of model mis-specification, may be derived from a different source.
@@ 2 The Ensemble Kalman Filter
Following from expository discussion of the Gaussian case in  Subsection 2.1, in Subsection 2.2 we define the specific version of the mean-field ensemble Kalman filter that we analyze here; other versions may be found in [8] and will be amenable to similar analyses. In Subsection 2.3 we prove our main stability theorem, showing that the error between the true filter and its Gaussian projection may be used to control the error between the true filter and the ensemble Kalman filter. In Subsection 2.4 we prove a corollary to this theorem, establishing that the mean field ensemble Kalman filter accurately approximates the true filter for a specific class of non-Gaussian problems.

### v6

# The Mean Field Ensemble Kalman Filter: Near-Gaussian Setting

@@ Abstract
The ensemble Kalman filter is widely used in applications because, for high dimensional filtering problems, it has a robustness that is not shared for example by the particle filter; in particular it does not suffer from weight collapse. However, there is no theory which quantifies its accuracy as an approximation of the true filtering distribution, except in the Gaussian setting. To address this issue we provide the first analysis of the accuracy of the ensemble Kalman filter beyond the Gaussian setting. We prove two types of results: the first type comprise a stability estimate controlling the error made by the ensemble Kalman filter in terms of the difference between the true filtering distribution and a nearby Gaussian; and the second type use this stability result to show that, in a neighbourhood of Gaussian problems, the ensemble Kalman filter makes a small error, in comparison with the true filtering distribution. Our analysis is developed for the mean field ensemble Kalman filter. We rewrite the update equations for this filter, and for the true filtering distribution, in terms of maps on probability measures. We introduce a weighted total variation metric to estimate the distance between the two filters and we prove various stability estimates for the maps defining the evolution of the two filters, in this metric. Using these stability estimates we prove results of the first and second types, in the weighted total variation metric. We also provide a generalization of these results to the Gaussian projected filter, which can be viewed as a mean field description of the unscented Kalman filter.
@@ 1.1 Context
This paper is concerned with the study of partially and noisily observed dynamical systems. Filtering refers to the sequential updating of the probability distribution of the state of the (possibly stochastic) dynamical system, given partial noisy observations [asch2016data, MR3363508, reich2015probabilistic]. The Kalman filter, introduced in 1960, provides an explicit solution to this problem in the setting of linear dynamical systems, linear observations and additive Gaussian noise [kalman1960new]; the desired probability distribution is Gaussian and the Kalman filter provides explicit update formulae for the mean and covariance. The extended Kalman filter is a linearization-based methodology which was developed in the 1960s and 1970s to apply to situations beyond the linear-Gaussian setting [jazwinski2007stochastic]. It is, however, not practical in high dimensions, because of the need to compute, and sequentially update, large covariance matrices [ghil]. To address this issue, the ensemble Kalman filter was introduced in 1994 [evensen1994sequential], using ensemble-based low-rank approximations of the covariances, and is a method well-adapted to high dimensional problems. The unscented Kalman filter, introduced in 1997, provides an alternative approach to nonlinear and non-Gaussian problems [julier1997new], using quadrature to approximate covariance matrices, and is well-adapted to problems of moderate dimension. The particle filter [doucet2001sequential] is a provably convergent methodology for approximating the filtering distribution [cricsan1999interacting, MR3375889]. However it does not scale well to high dimensional problems [MR2459233, snyder2008obstacles]; this motivates the increasing adoption of ensemble Kalman methods.
@@
Over the last two decades ensemble Kalman filters have found widespread use in the geophysical sciences, are starting to be used in other application domains, and have been developed as a general purpose tool for solving inverse problems; for reviews of such developments see [2022arXiv220911371C, evensen2009data, evensen2022data]. Despite widespread and growing adoption, theory quantifying the accuracy of the ensemble Kalman filter, in relation to the true filtering distribution, is limited. Two exceptions are the important contributions [le2009large, mandel2011convergence] which concern accuracy in the large particle limit in the setting where the underlying filtering problem is Gaussian. However, there is no proof that the ensemble Kalman methodology can accurately approximate the desired filtering distribution beyond the Gaussian setting. Indeed, in general the methodology does not correctly reproduce the filtering distribution; for examples and analysis in the setting of filtering and Bayesian inverse problems see [apte2008bayesian, le2009large] and [MR3400030] respectively. The aim of our work is to address this issue by proving accuracy of the ensemble Kalman filter beyond the Gaussian setting; specifically our analysis applies when the true filtering distribution is close to Gaussian, after appropriate lifting to the joint space of state and data. We perform the analysis for the mean field ensemble Kalman filter, focusing on quantifying the effect of the Gaussian approximation underlying the ensemble Kalman filter. We also study the Gaussian projected filter; this filter can be seen as a mean field version of the unscented Kalman filter. Both the ensemble and unscented Kalman mean field models are defined in [2022arXiv220911371C].
@@
Although our study of the accuracy of ensemble Kalman filters, beyond the Gaussian setting, is new, there exists a growing body of literature analyzing ensemble Kalman methods from different perspectives. In the context of long-time behaviour see the papers [kelly2014well, tong2016nonlinear, MR3784489, de2018long] which focus primarily on the accuracy of the filter in approximating the true trajectory over long time intervals; in contrast the papers [gottwald2013mechanism, kelly2015concrete] demonstrate a mechanism for filter divergence, an obstacle to obtaining accuracy. Localization is widely used in practice and important to consider in overall understanding of ensemble Kalman filter performance; see [tong2018performance, tong2023localized]. For analysis of filters in high dimensions see [snyder2015performance, majda2018performance]. For continuous-time and mean-field limits see [lange2020mean, lange2021continuous, ertel2022analysis]. In the context of inverse problems see [iglesias2013ensemble, schillings2017analysis, schillings2018convergence, huang2022iterated, huang2022efficient].
@@ 1.3 Notation
where $\mathrm{KL}(\mu\|\nu)$ is the Kullback–Leibler (KL) divergence of $\mu$ from $\nu$, defined in (B.30). Note that $\mathsf{G}$ defines a nonlinear mapping. We refer to $\mathsf{G}$ as a projection because of its characterization as finding closest point to $\mu$ with respect to the $\mathrm{KL}(\mu\|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ divergence. Throughout this paper all probability measures have continuous Lebesgue density, because of our assumptions concerning the noise structure in the dynamics model and the data acquisition model. Thus we abuse notation by using the same symbols for probability measures and their densities. For a probability measure $\mu\in\mathcal{P}(^{n})$, the notation $\mu(u)$ for $u\in^{n}$ refers to the Lebesgue density of $\mu$ evaluated at $u$, whereas $\mu[f]$ for a function $f\colon^{n}\to$ is a short-hand notation for $\int_{{}^{n}}f\mathrm{d}\mu$.
@@
Let $g\colon^{n}\to[1,\infty)$ be given by $g(v)=1+\lvert v\rvert^{2}$. We define the weighted total variation metric $d_{g}\colon\mathcal{P}^{2}(^{n})\times\mathcal{P}^{2}(^{n})\to$ by
@@
Metrics of this type have been employed previously in the literature. See for example the early reference [MR1174380], where a weighted total variation metric is used for stuyding ergodicity for Markov chains, and the reference [MR4074698], where a similar metric is employed in the context of perturbation of Markov chains. See also [MR2857021], where a direct proof of Harris’ ergodic theorem relying on appropriate weighted total variation norms is given, as well as [MR2459454, Theorem 6.15], which states that Wasserstein distances can be controlled by weighted total variation metrics. There are other dual metrics on probability measures commonly employed in the literature, such as the so-called maximum mean discrepancy, which is based on a dual formulation in a reproducing kernel Hilbert space; see e.g. [tolstikhin2016minimax].
@@
If $\mu_{1},\mu_{2}$ have Lebesgue densities $\rho_{1}$, $\rho_{2}$, then the weighted metric in Definition 1 satisfies
@@
Unlike the usual total variation distance, this weighted total variation metric enables control of the differences $\lvert\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\rvert$ and $\lVert\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\rVert$; several lemmas used to prove our main results rely on this control. More precisely, it is possible to show that for all $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ with finite second moments, it holds that
@@ 1.4.1 General Case
Here $\{u_{j}\}_{j\in\llbracket 0,J\rrbracket}$ is the unknown state, evolving in d, and $\{y_{j}\}_{j\in\llbracket 0,J\rrbracket}$ are the observations, evolving in K. We assume throughout the paper that the covariance matrices $\mathit{\Sigma},\mathit{\Gamma}$ are positive definite, that the function $h\colon^{d}\to^{K}$ is continuous, and that the initial state is distributed according to a Gaussian distribution $u_{0}\sim\mathcal{N}(m_{0},\mathit{C}_{0})$ and that the following independence assumption is satisfied:
@@
The objective of probabilistic filtering is to sequentially estimate the conditional distribution of the unknown state, given the data, as new data arrives. The true filtering distribution $\mu_{j}$ is the conditional distribution of the state $u_{j}$ given a realization $Y_{j}^{\dagger}=\{y_{1}^{\dagger},\dotsc,y_{j}^{\dagger}\}$ of the data process up to step $j$. Data $Y_{j}$ may be thought of as arising from a realization of (1.2); but the case of model mis-specification, where $Y_{j}$ does not necessarily arise from (1.2), is also of interest.
It is well-known [MR3363508, reich2015probabilistic] that the true filtering distribution evolves according to
@@
where $\mathsf{P}$ and $\mathsf{L}_{j}$ are maps on probability measures, referred to respectively as the prediction and analysis steps in the data assimilation community [asch2016data]. The operator $\mathsf{P}\colon\mathcal{P}(^{n})\to\mathcal{P}(^{n})$ is linear and defined by the Markov kernel associated with the stochastic dynamics (1.2a). Its action on a probability measure with Lebesgue density $\mu$ reads
@@
The operator $\mathsf{L}_{j}\colon\mathcal{P}(^{n})\to\mathcal{P}(^{n})$ is a nonlinear map which formalizes the incorporation of the new datum $y^{\dagger}_{j+1}$ using Bayes’ theorem. Its action on a probability measure in $\mathcal{P}(^{d})$ with Lebesgue density $\mu$ reads
@@
The operator $\mathsf{L}_{j}$ effects a reweighting of the measure to which it applies, with more weight assigned to the state values that are consistent with the observation. It is convenient in this work to decompose the analysis map $\mathsf{L}_{j}$ into the composition $\mathsf{B}_{j}\mathsf{Q}$. The action of the operator $\mathsf{Q}\colon\mathcal{P}(^{d})\to\mathcal{P}(^{d}\times^{K})$ on a probability measure in $\mathcal{P}(^{d})$ with Lebesgue density $\mu$ is given by
@@
On the other hand, the action of $\mathsf{B}_{j}\colon\mathcal{P}_{c}(^{d}\times^{K})\to\mathcal{P}(^{d})$ on a probability measure with continuous Lebesgue density $\mu$ is given by
@@
We note that for any $\mu\in\mathcal{P}(^{d})$, the measure $\mathsf{Q}\mathsf{P}\mu\in\mathcal{P}_{c}(^{d}\times^{K})$ has continuous Lebesgue density, and so the iteration (1.9) is well defined.
@@ 1.4.2 Gaussian Case
Now consider the setting in which $\Psi(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})=M\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}$, and $h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})=H\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}$ for matrices $\mathit{M}\in^{d\times d}$ and $\mathit{H}\in^{K\times d}.$ In this case the filtering distribution is Gaussian: $\mu_{j}=\mathcal{N}(m_{j},C_{j}).$ The Kalman filter [kalman1960new] gives explicit evolution equations for the pair $(m_{j},C_{j})\in^{d}\times^{d\times d}.$ To write these down it is helpful to note that $\widehat{\mu}_{j+1}:={\mathsf{P}}\mu_{j}$ is also Gaussian: $\widehat{\mu}_{j+1}=\mathcal{N}(\widehat{m}_{j},\widehat{\mathit{C}}_{j}).$
@@
where $\{{y}^{\dagger}_{j}\}$ is the observation. Recall that this observation may be thought of as arising from a realization of (1.2); but, in the case of model mis-specification, may be derived from a different source.
@@ 2 The Ensemble Kalman Filter
Following from expository discussion of the Gaussian case in  Subsection 2.1, in Subsection 2.2 we define the specific version of the mean-field ensemble Kalman filter that we analyze here; other versions may be found in [2022arXiv220911371C] and will be amenable to similar analyses. In Subsection 2.3 we prove our main stability theorem, showing that the error between the true filter and its Gaussian projection may be used to control the error between the true filter and the ensemble Kalman filter. In Subsection 2.4 we prove a corollary to this theorem, establishing that the mean field ensemble Kalman filter accurately approximates the true filter for a specific class of non-Gaussian problems.